OpenStudy (anonymous):

Solve the equation. -2-2+3x=6 can someone show me step by step how to do this...

OpenStudy (angela210793):

-4+3x=6 add 4 both sides 3x=10 divide by 3 both ides x=10/3

OpenStudy (anonymous):

-4+3x=6 -4+3x+4=6+4 3x=10 3x/3=10/3 x=10/3

OpenStudy (anonymous):

-2-2+3x=6 first subtract the 2 from the -2: -2-2=-4 so -4+3x=6 now add 4 to each side -4+4+3x=6+4 simplify 3x=10 divide each side by 3 3x/3=10/3 simplify x=10/3

OpenStudy (anonymous):

-2-2+3x=6 -4+3x=6 3x=6 +4 x=10/3